OpenStudy (anonymous):

Which of the following items has the lowest unit price? 4 for $5.00 $1.22 each 6 for $7.44 3 for $3.60

OpenStudy (btaylor):

A 5/4 = 1.25 each B = 1.22 each C 7.44/6 = 1.24 each ----> D 3.60/3 = 1.20 each

OpenStudy (anonymous):

thanks